OpenStudy (anonymous):

Please help me... the time required to do a job varies inversely as the number of people working. It takes 6hrs for 3 bricklayers to build a park wall. How long will it take 7 bricklayers to complete the job?

OpenStudy (goldphenoix):

You can use a proportion to solve this:  I hope this help! :)

OpenStudy (anonymous):

it does thanks

OpenStudy (radar):

That is strange, as 3 bricklayers can build it in 6 hours, you would expect with more bricklayers it would take less than 6 hrs, not more like 14 hrs. remember it is an inverse relationship.

OpenStudy (goldphenoix):

Hmm. You do got a point.

OpenStudy (goldphenoix):

Oh no, so sorry, marigarcia. T~T I should've checked my work.

OpenStudy (anonymous):

ok Radar so how is the equation will look like..

OpenStudy (anonymous):

goldPhenoix its ok..

OpenStudy (radar):

I am hoping GoldPhenoix will show you how. I know if there was 6 bricklayers they should build the wall in 3 hrs.

OpenStudy (anonymous):

yeah that's what I'm thinking.. so Radar you are not going to help me???

OpenStudy (radar):

I am trying to come up with a inverse relationship, I'm thinking, give me some time to figure this thing out.

OpenStudy (anonymous):

radar ok.. thanks

OpenStudy (goldphenoix):

I'm thinking of doing 3/7 and then multiply it by 6. Would that give you the answer?

OpenStudy (anonymous):

GoldPhenoix so how the equation will be?/

OpenStudy (radar):

An inverse relationship would be something like y= k/x where k is the constant of proportion, x is the variable, and y is the result. I think GoldPhenois is on the right track

OpenStudy (radar):

6=k/3 so k=18 x=18/7 seems that would be the answer. I am rusty at this.

OpenStudy (anonymous):

Radar that's the same equation that I have.

OpenStudy (radar):

And GoldPhenoix comes out the same, I think that is a reasonable answer let me try it with the 6 bricklayers y=18/6 =3 hrs. It makes sense.

OpenStudy (anonymous):

the answer is 2hrs and 4/7

OpenStudy (radar):

Do you get instant feedback when you input an answer, if so, was that correct?

OpenStudy (anonymous):

I was confused but when you explain it make more sense..

OpenStudy (anonymous):

know can you help me with this problem because I'm not good at solving story problems

OpenStudy (radar):

I believe GoldPhenoix first post is true when their is a direct relationship. In this case it was an inverse relationship. His later post took that into consideration. I am going to award a medal for him, you deserve one also, because you also stuck with it and see how it is done.

OpenStudy (anonymous):

thanks for helping me and for the award..

OpenStudy (radar):

The photo for your avatar is that a doll or is it your child?

OpenStudy (radar):

It only allowed me to make one award. But you deserve an award.

OpenStudy (anonymous):

would you help me to solve this.... A student traveled 244 km in 17 days. At the same ratio, how far would the student travel in 85 days.

OpenStudy (anonymous):

radar its ok with the reward gave it to Goldphenoix.. about the photo its my nice..

OpenStudy (anonymous):

I think the equation for this one will be something like this: r=d/t, I'm I right

OpenStudy (radar):

That one seems easier. First find out what is travel is for one day by dividing 244 by 17 oops they want you to use ratios. 17:244 :: 85:X